# Answers to frequently asked questions about the pulsar timing array Hellings and Downs curve

### v2

@@ 4 Answers to frequently asked questions
Consider how the $N(N-1)$ terms of the second sum add together. If the phases $\phi_{n}$ of the $N$ sources are random and independent, then the cosine has values that lie in the interval $[-1,1]$, so the sum includes both positive and negative terms. If the number of GW sources $N$ is large, then this second sum can be thought of as doing a random walk with $N(N-1)$ steps to the left and right along the real number line. It thus contributes a value whose typical magnitude is of order $\sqrt{N(N-1))}\approx N$, which is comparable to the magnitude of the first term. So, even after averaging over many pulsars uniformly distributed over the sky, the second term gives rise to a deviation away from the Hellings and Downs curve. The mean-squared deviation is exactly the cosmic variance.
@@
We can easily estimate the size of the cosmic variance, starting from (4). The mean value, assuming independent random phases, and averaging over many different universes, is the first term in (4). The deviation away from that (unobservable) mean, in any given realization of the universe, is given by the second term in (4). If the GW sources are uniformly distributed over the two-sphere, then the distribution of the angles $\beta_{nm}$ is proportional to $\sin\beta_{nm}$. Thus, the expected squared deviation of the pulsar-averaged correlation away from the Hellings and Downs curve, as a function of $\gamma$, has the form
$\sigma^{2}_{\rm cosmic}(\gamma)\propto\int_{0}^{\pi}{\rm d}\beta\>\sin\beta\,\mu^{2}(\gamma,\beta)\,.$ (1cdmatbebgbm)
@@
For a more realistic set of GW sources, the full formula is derived in [9, 19, 21]. The basic ideas are the same, so it is not surprising that the equation is identical to (1cdmatbebgbm), except that it also includes the overall coefficient, which we have omitted. In particular, the angular dependence on $\gamma$ is exactly as given in (1cdmatbebgbm), with an overall coefficient that correctly weights the contributions of GW sources radiating at different frequencies.
@@
    Assuming noise-free observations, are the fluctuations away from the expected Hellings and Downs curve due solely to the pulsar-term contributions to the timing residual response?
@@
Answer: While the match for any single angular-separation bin is only at the 1-$\sigma$ level (as it should be for 1-$\sigma$ error bars), the significance comes from the agreement between the measured and expected correlations combined over all pulsar-pair correlations. In other words, one needs to compute a single “signal-to-noise ratio” by summing inner products of the observed pulsar-pair correlations with the expected Hellings and Downs “waveform”. (This is similar to what LIGO does when it compares snippets of its noisy time-domain data to noise-free chirp templates.) The individual terms of this sum are inversely weighted by estimates of their expected squared values in the absence of the Hellings and Downs correlations. So, even though the match at single angular separation is not very close, the overall agreement is quite good.
In fact, NANOGrav obtains significance estimates from two types of analyses:
@@
(i) A $3.8\sigma$ significance determined using frequentist detection statistics, such as the noise-weighted inner product of the observed correlations with the expected Hellings and Downs correlation, as described above. This is the “stochastic equivalent” of the deterministic matched filter statistic used by LIGO to search for individual compact binary coalescence “chirp” signals. The distribution that would be expected in the absence of a GWB is estimated by artificially introducing phase shifts into the timing residual data for each pulsar, which removes the Hellings and Downs correlations (see the right panel of Figure 3 in [4]). The observed value of the detection statistic lies in the “tail” of the null distribution, and thus corresponds to high significance.
(ii) A $3.0\sigma$ significance obtained via a Bayesian analysis. This analysis compares the evidence for a data model containing the Hellings and Downs spatial correlations to a model without these correlations. The ratio of the evidences is 226, which is the “betting odds” for which of the two models is correct. The significance of the value 226 is then estimated from the null distribution of the evidence ratio, again constructed using phase shifts (see both Figure 2 and the left panel of Figure 3 in [4]).
@@
So, what is shown in Figure 1 for the NANOGrav 15-year analysis [4] should be thought of as a “self-consistency check” of the pulsar-pair cross-correlations, calculated under the hypothesis that a GWB with Hellings and Downs correlations matching the Hellings and Downs curve is present in the data, with an overall amplitude estimated from auto-correlation measurements, and assuming that there are no unknown sources of noise with similar spatial correlations. The estimated correlations and error bars are optimally weighted combinations of approximately 150 correlation measurements in each angular separation bin, which include the covariances between the correlations induced by the GWB. If the observed data are consistent with the Hellings and Downs correlation model, then these point estimates and 1-$\sigma$ error bars will cross the Hellings and Downs curve approximately 70% of the time on average. See [19] for a more detailed discussion.
@@ 5 Summary
We calculated the effect of gravitational waves on the travel time of light, and showed how the Hellings and Downs curve arises when correlating timing residuals from PTA pulsars. We then answered a number of frequently asked questions (FAQs) about that curve.
@@
This paper and its “FAQ sheet” format arose from discussions that JDR had with his wife and with several students attending the “MaNiTou Summer School on GWs,” held from 3 July to 8 July 2023, in Nice, France. JDR acknowledges past conversations with numerous NANOGrav colleagues about various aspects of these topics, and financial support from the NSF Physics Frontier Center Award PFC-2020265. He also thanks colleagues Nelson Christensen and Marie-Anne Bizouard for their support and hospitality while visiting the Artemis Lab, Observatoire Cote D’Azur in June and July 2023, where the initial drafts of this paper were written.

### v3

@@ 4 Answers to frequently asked questions
It is easy to quantify the amount by which a typical universe will vary away from the average of many universes, which is the HD curve. Consider how the $N(N-1)$ terms of the second sum add together. If the phases $\phi_{n}$ of the $N$ sources are random and independent, then the cosine has values that lie in the interval $[-1,1]$, so the sum includes both positive and negative terms. If the number of GW sources $N$ is large, then this second sum can be thought of as doing a random walk with $N(N-1)$ steps to the left and right along the real number line. It thus contributes a value whose typical magnitude is of order $\sqrt{N(N-1))}\approx N$, which is comparable to the magnitude of the first term. So, even if the number of sources $N\to\infty$, the second term gives rise to a nonzero deviation away from the Hellings and Downs curve. The mean-squared value of this deviation is exactly the cosmic variance.
@@
We can easily estimate the size of the cosmic variance, starting from (4). The mean value, assuming independent random phases, and averaging over many different universes, is the first term in (4). The deviation away from that (unobservable) mean, in any given realization of the universe, is given by the second term in (4). If the GW sources are uniformly distributed over the two-sphere, then the distribution of the angles $\beta_{nm}$ is proportional to $\sin\beta_{nm}$. Thus, the expected squared deviation of the pulsar-averaged correlation away from the Hellings and Downs curve, as a function of $\gamma$, has the form [^10] The $\chi^{\phantom{{}^{\prime}}}_{nm}$ do not affect the distribution over $\beta$ in (1cdmaubfbhbp): they are eliminated from the second moment when $\cos\Delta_{nm}\cos\Delta_{kl}$ for $n\neq m$ and $k\neq l$ is averaged over the random phases $\phi_{j}$. To see this, write the cosines as exponentials. The average has four terms. A typical term is ${\rm e}^{i(\Delta_{nm}+\Delta_{kl})}$, whose random phase average is $\delta_{nl}\delta_{mk}{\rm e}^{-2i(\chi^{\phantom{{}^{\prime}}}_{nm}+\chi^{\phantom{{}^{\prime}}}_{kl})}$. The $\chi$’s cancel and drop out because $\chi^{\phantom{{}^{\prime}}}_{nm}+\chi^{\phantom{{}^{\prime}}}_{mn}$ vanishes.
$\sigma^{2}_{\rm cosmic}(\gamma)\propto\int_{0}^{\pi}{\rm d}\beta\>\sin\beta\,\mu^{2}(\gamma,\beta)\,\,\,=2\sum_{l=2}^{\infty}(2l+1)c^{2}_{l}P^{2}_{l}(\cos\gamma)\,.$ (1cdmaubfbhbp)
@@
In the final equality, $c_{l}$ are defined by (1cdmaubd); see [17, Eq. (8.3)]. For a more realistic set of GW sources, the full formula is derived in [5, 23, 25]. The basic ideas are the same, so it is not surprising that the equation is identical to (1cdmaubfbhbp), except that it also includes the overall coefficient, which we have omitted. In particular, the angular dependence on $\gamma$ is exactly as given in (1cdmaubfbhbp), with an overall coefficient that correctly weights the contributions of GW sources radiating at different frequencies.
@@
    Assuming noise-free observations, are the fluctuations away from the expected Hellings and Downs curve only due to the pulsar-term contributions to the timing residual response?
@@
Answer: While the match for any single angular-separation bin is only at the 1-$\sigma$ level (as it should be for 1-$\sigma$ error bars), the significance comes from the agreement between the measured and expected correlations combined over all pulsar-pair correlations. In other words, one needs to compute a single “signal-to-noise ratio” by summing inner products of the observed pulsar-pair correlations with the expected Hellings and Downs “waveform”. (This is similar to what LIGO does when it compares snippets of its noisy time-domain data to noise-free chirp templates.) The individual terms of this sum are inversely weighted by estimates of their expected squared values in the absence of the Hellings and Downs correlations. So, even though the match at the individual angular separations are not very close, the overall agreement is quite good.
In fact, NANOGrav obtains significance estimates [^11] Although we focus on NANOGrav, the general discussion applies to all PTAs apart from CPTA, which uses a different detection statistic. from two types of analyses:
@@
(i) A $3.8\sigma$ significance determined using frequentist detection statistics, such as the noise-weighted inner product of the observed correlations with the expected Hellings and Downs correlation, as described above [26]. This is the “stochastic equivalent” of the deterministic matched filter statistic used by LIGO to search for individual compact binary coalescence “chirp” signals. The distribution that would be expected in the absence of a GWB is estimated by artificially introducing phase shifts into the timing residual data for each pulsar, which removes the Hellings and Downs correlations (see the right panel of Figure 3 in [4]). The observed value of the detection statistic lies in the “tail” of the null distribution, and thus corresponds to high significance.
@@
(ii) A $3.0\sigma$ significance obtained via a Bayesian analysis. This analysis compares the evidence for a data model containing the Hellings and Downs spatial correlations to a model without these correlations. The ratio of the evidences is 226, which is the “betting odds” for which of the two models is correct. The significance of the value 226 is then estimated from the null distribution of the evidence ratio, again constructed using phase shifts (see both Figure 2 and the left panel of Figure 3 in [4]). So, what is shown in Figure 1 for the NANOGrav 15-year analysis [4] should be thought of as a “self-consistency check” of the pulsar-pair cross-correlations, calculated under the hypothesis that a GWB with Hellings and Downs correlations matching the Hellings and Downs curve is present in the data, with an overall amplitude estimated from auto-correlation measurements, and assuming that there are no unknown sources of noise with similar spatial correlations. The estimated correlations and error bars are optimally weighted combinations of approximately 150 correlation measurements in each angular separation bin, taking into account the covariances between the correlations induced by the GWB. If the observed data are consistent with the Hellings and Downs correlation model, then these point estimates and 1-$\sigma$ error bars will cross the Hellings and Downs curve approximately 68% of the time on average. See [23] for a more detailed discussion.
  • Q10:
    If deviations from the Hellings and Downs curve are detected, what does that mean?
Answer: The reconstruction of the Hellings and Downs correlation is necessarily imperfect, for several reasons. First, because in practice, the number of pulsar pairs is finite, whereas the Hellings and Downs curve is obtained by averaging over an infinite number of pulsar pairs separated by angle $\gamma$. Second, because pulsars are not perfect clocks, and their pulse arrival times are modified by effects that cannot be completely modeled and removed [27, Sec. 4]. However, if the number of pulsar pairs is large enough [23, Fig. 6] and if the duration of the observations is long enough, then both of these issues can be overcome.
So, what if deviations from the Hellings and Downs curve are found, when there is enough high-quality data from enough pulsars, so that neither of these practical issues can be responsible? In this case, we have to ask about the size of the deviations, and about how they are correlated as a function of the angle $\gamma$. If the deviations are consistent with the predictions of cosmic variance and covariance [5, 23], it suggests that the GWs arise from a large number of interfering sources, consistent with a central-limit-theorem type of process. Indeed, an exact match to the Hellings and Downs curve would suggest the contrary: that the GWs are being produced by a small number of GW sources, radiating at distinct non-interfering frequencies. See [23, Sec. VIII] and [5, Conclusion] for further discussion of this point.
@@ 5 Summary
We calculated the effect of GWs on the travel time of electromagnetic signals such as light or radio waves, and showed how the Hellings and Downs curve arises when correlating timing residuals from PTA pulsars. We then answered a number of frequently asked questions (FAQs) about that curve.
@@
This paper and its “FAQ sheet” format arose from discussions that JDR had with his wife and with several students attending the “MaNiTou Summer School on GWs,” held from 3 July to 8 July 2023, in Nice, France. JDR acknowledges past conversations with numerous NANOGrav colleagues about various aspects of these topics, and financial support from the NSF Physics Frontier Center Award PFC-2020265. He also thanks colleagues Nelson Christensen and Marie-Anne Bizouard for their support and hospitality while visiting the Artemis Lab, Observatoire Cote D’Azur in June and July 2023, where the initial drafts of this paper were written. The authors thank the referees for helpful comments and suggestions, in particular, regarding the close analogy between the short and long-arm antenna patterns for GW and electromagnetic detectors / antennae.